# Subvarieties of pointed Abelian $\ell$-groups

### v1

@@ Abstract
This paper provides a complete classification of the subvarieties and subquasivarieties of pointed Abelian lattice-ordered groups ($\ell$-groups) that are generated by chains. We present two complementary approaches to achieve this classification.
First, using purely $\ell$-group-theoretic methods, we analyze the structure of lexicographic products and radicals to identify all join-irreducible members of the lattice of subvarieties of positively pointed $\ell$-groups. We provide a novel equational basis for each of these subvarieties, leading to a complete description of the entire subvariety lattice. As a direct application, our $\ell$-group-theoretic classification yields an alternative, self-contained proof of Komori’s celebrated classification of subvarieties of MV-algebras.
Second, we explore the connection to MV-algebras via Mundici’s $\Gamma$ functor. We prove that this functor preserves universal classes, a result of independent model-theoretic interest. This allows us to lift the classification of universal classes of MV-chains, due to Gispert, to a complete classification of universal classes of totally ordered pointed Abelian $\ell$-groups. As a direct consequence, we obtain a full description of the corresponding lattice of subquasivarieties. These results offer a comprehensive structural understanding of one of the most fundamental classes of ordered algebraic structures.
Keywords: varieties, pointed Abelian $\ell$-groups, quasivarieties, universal classes, lexicographic product, axiomatization, Abelian logic, MV-algebras, Mundici functor, Komori classification
@@ 1 Introduction
Lattice-ordered groups ($\ell$-groups) represent a foundational class of algebraic structures with a rich history of celebrated results [18, 16, 15, 5, 11]. The field remains an active area of research, largely because the lattice of subvarieties of $\ell$-groups is notoriously complex and not yet fully understood [22]. Even the simplest case, the variety of Abelian $\ell$-groups, presents a deep mathematical challenge.
Another extremely important class of algebraic structures is that of MV-algebras. Originally introduced as the algebraic semantics for Łukasiewicz logic [29], these algebras have been the subject of extensive study for decades [7, 13, 23, 25]. One of the most essential tools in their study is the celebrated Mundici functor, a categorical equivalence between MV-algebras and Abelian $\ell$-groups with a strong unit. This functor provides a powerful bridge between the two fields, allowing many fundamental results in the theory of MV-algebras to be obtained by translating problems into the more established context of $\ell$-groups.
However, the theory of MV-algebras is in some respects richer than that of Abelian $\ell$-groups, primarily because its language contains an additional designated constant. A similar level of expressive power can be achieved in the theory of Abelian $\ell$-groups by augmenting their language with an additional constant, which gives rise to the class of pointed Abelian $\ell$-groups. As this article demonstrates, this class allows one to reason about MV-algebras while remaining entirely within the realm of Abelian $\ell$-groups. Furthermore, this class forms a variety, making it a natural and well-behaved extension of the class of Abelian $\ell$-groups with a strong unit, which is not even first-order definable.
@@
    Second, we develop our theory semantically from first principles within the theory of Abelian $\ell$-groups, rather than relying on the Mundici functor and the theory of MV-algebras. This contrasts with the more syntactic approach of [28] and allows us to prove more general theorems (e.g., compare our Theorem 3.10 with [28, Lemma 4.6]). As a final consequence of our framework, we derive the original Komori classification in Theorem 6.3.
@@
    Finally, we will use Gispert’s classification of universal classes of MV-chains and quasivarieties of MV-algebras generated by chains from [14] to provide the corresponding classifications for pointed Abelian $\ell$-groups.
The paper is structured as follows. Section 2 introduces the necessary preliminaries from universal algebra and the theory of Abelian $\ell$-groups. Section 3 develops the core technical results concerning lexicographic decompositions of finitely generated totally ordered $\ell$-groups, culminating in Theorem 3.10. Section 4 provides a semantic characterization of the varieties of Abelian $\ell$-groups, presenting novel and more elegant proofs for some known results (e.g., Lemma 4.8). The central contribution of the paper is presented in Section 5, which gives a complete classification and axiomatization of all subvarieties of pointed Abelian $\ell$-groups. Finally, Section 6 applies our results to the theory of MV-algebras, providing an alternative proof of Komori’s classification. Moreover, by using Gispert’s classification of universal classes of MV-chains, we provide a similar classification for $\ell$-group chains and consequently obtain a classification of all quasivarieties generated by $\ell$-group chains.
@@ 2 Preliminaries
In this section we introduce some basic terminology and well-known results about Abelian $\ell$-groups. We assume that the reader has a basic background in universal algebra. No advanced background is assumed throughout the paper. We will use the standard notations $\mathbf{H},\mathbf{I},\mathbf{S},\mathbf{P},\mathbf{P}_{\mathrm{U}}$ to denote closure under homomorphisms, isomorphisms, subalgebras, products, and ultrapowers, respectively. We will write ${\boldsymbol{\mathit{A}}}\in{{\mathbb{K}}}_{SI}$ iff the algebra $\boldsymbol{\mathit{A}}$} is a subdirectly irreducible element of the class $\mathbb{K}$}}.
@@
    ${{\mathbb{K}}}=\mathbf{ISP}({{\mathbb{K}}}_{SI})$.
  3. 3.
    ${{\mathbb{K}}}=\mathbf{ISPP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{A}}}\in{{\mathbb{K}}}_{SI}\mid{\boldsymbol{\mathit{A}}}\text{ is finitely generated}\})$.
  4. 4.
@@
  1. 5.
    $\mathbb{K}$}} is the class of all models of some theory, all of whose axioms are equations with the right side equal to 0.
@@
Let ${{\mathbb{K}}}\cup\{{\boldsymbol{\mathit{A}}}\}$ be a class of algebras of the same signature. Let $F$ be a finite subset of $A$ and ${\boldsymbol{\mathit{B}}}\in{{\mathbb{K}}}$. We say that a mapping $f_{F}:F\rightarrow B$ is a partial embedding of the set $F$ from $\boldsymbol{\mathit{A}}$} into $\boldsymbol{\mathit{B}}$} if $f_{F}$ is a one-to-one mapping such that for each $a_{1},\dots,a_{n}\in F$ and for each $\lambda\in\mathcal{L}$ such that $\lambda^{{\boldsymbol{\mathit{A}}}}(a_{1},\dots,a_{n})\in F$ we have
@@
We say that $\boldsymbol{\mathit{A}}$} is partially embeddable into the class $\mathbb{K}$}} if for each finite set $F\subseteq A$ there is a ${\boldsymbol{\mathit{B}}}\in{{\mathbb{K}}}$ and there is a partial embedding $f_{F}$ of $F$ into $\boldsymbol{\mathit{B}}$}. If $\boldsymbol{\mathit{A}}$} is partially embeddable into $\{{\boldsymbol{\mathit{B}}}\}$ we just say $\boldsymbol{\mathit{A}}$} is partially embeddable into $\boldsymbol{\mathit{B}}$}.
Lemma 2.3.
@@
    $\mathbb{K}$}} is closed under partial embeddings (i.e. for any algebra $\boldsymbol{\mathit{A}}$} that is partially embeddable into $\mathbb{K}$}} we get ${\boldsymbol{\mathit{A}}}\in{{\mathbb{K}}}$).
@@
An algebra ${\boldsymbol{\mathit{A}}}=\langle{A,+,-,\lor,\land,0}\rangle$ is an Abelian $\ell$-group if $\langle{A,+,-,0}\rangle$ is an Abelian group, $\langle{A,\lor,\land}\rangle$ is a lattice and $\boldsymbol{\mathit{A}}$} satisfies the monotonicity condition, that is, for each $x,y,z\in A$ we get $x\leq y$ implies $x+z\leq y+z$.
It is well-known (see, e.g. [12, Chapter V]), that the defining conditions of Abelian $\ell$-groups can be expressed by means of equations, so they form a variety which we denote by $\mathbb{AL}$. Also, it is well-known that all Abelian $\ell$-groups are torsion-free.
We will denote by $\boldsymbol{\mathit{Z}}$} and $\boldsymbol{\mathit{R}}$} the $\ell$-groups of integers and reals with the underlying universes $\mathbb{Z}$ and $\mathbb{R}$.
@@
An Abelian $\ell$-group is subdirectly irreducible iff it is totally ordered.
@@
All totally ordered Abelian $\ell$-groups are universally equivalent. Equivalently, for every totally ordered Abelian $\ell$-group $\boldsymbol{\mathit{A}}$} we have $\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{A}}})=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}})$.
From Theorems 2.5 and 2.6 one can easily derive the following:
@@
The following are equivalent for any Abelian $\ell$-group $\boldsymbol{\mathit{A}}$}:
@@
Let $\boldsymbol{\mathit{A}}$} be an $\ell$-subgroup of $\boldsymbol{\mathit{R}}$}. Then $A$ is a dense subset of $\mathbb{R}$ (with respect to the standard topology on $\mathbb{R}$) or ${\boldsymbol{\mathit{A}}}\cong{\boldsymbol{\mathit{Z}}}$.
@@
Since multiplication by any element of $\mathbb{R}$ is an $\ell$-group automorphism on $\boldsymbol{\mathit{R}}$} we can without loss of generality assume $1\in{\boldsymbol{\mathit{A}}}$. If $\boldsymbol{\mathit{A}}$} contains ${\boldsymbol{\mathit{Z}}}[\xi]$ for some $\xi\in\mathbb{R}\setminus\mathbb{Q}$, it is easy to show (using Euclidean algorithm on $1$ and $\xi$) that $A$ must be a dense subset of $\mathbb{R}$.
Otherwise, ${\boldsymbol{\mathit{A}}}\in\mathbf{S}({\boldsymbol{\mathit{Q}}})$. By [1, Theorem 2] and [1, Corollary 2] for every non-cyclic ${\boldsymbol{\mathit{A}}}\in\mathbf{S}({\boldsymbol{\mathit{Q}}})$ for each $\epsilon\in\mathbb{R}$ there is $a\in A$ such that $0<a\leq\epsilon$. Since $\boldsymbol{\mathit{A}}$} is also closed under addition, AA has to be a dense subset of $\mathbb{R}$. If $\boldsymbol{\mathit{A}}$} is cyclic then ${\boldsymbol{\mathit{A}}}\cong{\boldsymbol{\mathit{Z}}}$. ∎
@@
It is well known that congruences on Abelian $\ell$-groups coincide with convex subgroups, see e.g. [12].
Despite the fact that the multiplication symbol $\cdot$ is not present in the language of Abelian $\ell$-groups we will commonly use it in the traditional meaning of the iterated addition. Clearly, the classes $p\mathbb{AL},p\mathbb{AL}^{+},p\mathbb{AL}^{-}$ and $p\mathbb{AL}^{0}$ are varieties. Let us note that for an Abelian $\ell$-group $\boldsymbol{\mathit{A}}$} and $a\in A$ it holds that $\boldsymbol{\mathit{A}}$} is subdirectly irreducible iff ${\boldsymbol{\mathit{A}}}_{a}$ is subdirectly irreducible. Therefore, using Theorem 2.5 we obtain that a pointed Abelian $\ell$-group is subdirectly irreducible iff it is totally ordered. Another important fact is that the variety $p\mathbb{AL}^{0}$ is obviously term equivalent to $\mathbb{AL}$. The following can be proved.
@@
    $p\mathbb{AL}_{SI}=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$.
@@
    $p\mathbb{AL}_{SI}^{+}=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$.
@@
    $p\mathbb{AL}_{SI}^{-}=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0})$.
@@
    $p\mathbb{AL}_{SI}^{0}=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{0})=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{0})$.
@@
Let us recall that pointed Abelian $\ell$-groups are subdirectly irreducible iff they are linearly ordered.
All the inclusions $\supseteq$ are trivial. We prove the remaining inclusions in a convenient order.
@@
    Follows, since the class $p\mathbb{AL}_{SI}^{0}$ is term equivalent to $\mathbb{AL}_{SI}.$
@@
    Let us fix an arbitrary ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}_{SI}$. Using Theorem 2.6 and [8, Lemma 4.3] we obtain ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{R}}}_{b}\mid b\in\mathbb{R}\})$. Since for $b\in\mathbb{R}$ we have ${\boldsymbol{\mathit{R}}}_{b}\cong{\boldsymbol{\mathit{R}}}_{-1}$, ${\boldsymbol{\mathit{R}}}_{b}\cong{\boldsymbol{\mathit{R}}}_{0}$, or ${\boldsymbol{\mathit{R}}}_{b}\cong{\boldsymbol{\mathit{R}}}_{1}$ we obtain ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$.
@@
    For ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}_{SI}^{0}$ we already know ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$. Therefore, we fix arbitrary ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}_{SI}^{+}\setminus p\mathbb{AL}_{SI}^{0}$. By the previous point we know that ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$. By [2, Theorem 5.6] we have
@@
    Since ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}_{SI}^{+}\setminus p\mathbb{AL}_{SI}^{0}$, it follows $a>0$. Therefore, ${\boldsymbol{\mathit{A}}}_{a}\notin\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1})$ since ${\boldsymbol{\mathit{A}}}_{a}\nvDash\mathtt{f}\leq 0$. Thus ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$.

### v2

@@ Abstract
This paper provides a complete classification of the subvarieties and subquasivarieties of pointed Abelian lattice-ordered groups ($\ell$-groups) that are generated by their totally ordered members. We present two complementary approaches to achieve this classification.
First, using purely $\ell$-group-theoretic methods, we analyze the structure of lexicographic products and values to identify all join-irreducible members of the lattice of subvarieties of positively pointed Abelian $\ell$-groups. We provide a novel equational basis for each of these subvarieties, leading to a complete description of the entire subvariety lattice. As a direct application, our $\ell$-group-theoretic classification yields an alternative, self-contained proof of Komori’s classification of subvarieties of MV-algebras.
Second, we explore the connection to MV-algebras via Mundici’s $\Gamma$ functor. We prove that this functor preserves universal classes, a result of independent model-theoretic interest. This allows us to lift the classification of universal classes of totally ordered MV-algebras, due to Gispert, to a complete classification of universal classes of totally ordered pointed Abelian $\ell$-groups. As a direct consequence, we obtain a full description of the corresponding lattice of subquasivarieties.
Keywords: varieties, pointed Abelian $\ell$-groups, quasivarieties, universal classes, lexicographic product, axiomatization, Abelian logic, MV-algebras, Mundici functor, Komori’s classification
@@ 1 Introduction
Lattice-ordered groups ($\ell$-groups) represent a foundational class of algebraic structures with a rich history of celebrated results [20, 18, 17, 6, 13]. The field remains an active area of research, largely because the lattice of subvarieties of $\ell$-groups is notoriously complex and not yet fully understood [25]. Even the simplest non-trivial case, the variety of Abelian $\ell$-groups, presents a deep mathematical challenge.
MV-algebras form another extremely important class of algebraic structures. Originally introduced as the algebraic semantics for Łukasiewicz logic [33], these algebras have been the subject of extensive study for decades [8, 15, 26, 29]. One of the most essential tools in their study is the celebrated Mundici functor, a categorical equivalence between MV-algebras and Abelian $\ell$-groups with a strong unit. This functor provides a powerful bridge between the two fields, allowing many fundamental results in the theory of MV-algebras to be obtained by translating problems into the more established context of $\ell$-groups.
However, the theory of MV-algebras is in some respects richer than that of Abelian $\ell$-groups, primarily because its language contains an additional designated constant. A similar level of expressive power can be achieved in the theory of Abelian $\ell$-groups by augmenting their language with an additional constant, which gives rise to the class of pointed Abelian $\ell$-groups. As this article demonstrates, this class allows one to reason about MV-algebras while remaining entirely within the realm of Abelian $\ell$-groups. Furthermore, this class forms a variety, making it a natural and well-behaved algebraic framework. This contrasts with the class of Abelian $\ell$-groups with a strong unit, which is not even first-order definable.
@@
    Second, we develop our theory of varieties semantically from first principles within the theory of Abelian $\ell$-groups, rather than relying on the Mundici functor and the theory of MV-algebras. This contrasts with the more syntactic approach of [32] and allows us to prove more general theorems (e.g., compare our Theorem 3.10 with [32, Lemma 4.6]). Moreover, this strategy provides us with more structural insight, and we believe readers will find our reasoning more structurally transparent than that presented in [32]. As a consequence of our framework, we derive Komori’s classification of subvarieties of MV-algebras in Theorem 6.3.
@@
    Finally, we will use Gispert’s classification of universal classes of totally ordered MV-algebras and quasivarieties of MV-algebras generated by their totally ordered members from [16] to provide the corresponding classification for pointed Abelian $\ell$-groups.
The paper is structured as follows. Section 2 introduces the necessary preliminaries from universal algebra and the theory of Abelian $\ell$-groups. Section 3 develops the core technical results concerning lexicographic decompositions of finitely generated totally ordered $\ell$-groups, culminating in Corollary 3.6 and Theorem 3.10. Section 4 provides a semantic characterization of the varieties of pointed Abelian $\ell$-groups, presenting novel and more elegant proofs for some known results (e.g., Lemma 4.8). The central contribution of the paper is presented in Section 5, which gives a complete classification and axiomatization of all subvarieties of pointed Abelian $\ell$-groups. Finally, Section 6 applies our results to the theory of MV-algebras, providing an alternative proof of Komori’s classification. Moreover, by using Gispert’s classification of universal classes of totally ordered MV-algebras, we provide a similar classification for totally ordered pointed Abelian $\ell$-groups and consequently obtain a classification of all quasivarieties generated by totally ordered pointed Abelian $\ell$-groups.
@@ 2 Preliminaries
In this section, we introduce some basic terminology and well-known results about Abelian $\ell$-groups. We assume the reader has a basic background in universal algebra, but no advanced prerequisites are required to follow the rest of the paper. We will use the standard notations $\mathbf{H,I,S,P,\mathbf{P}_{\mathrm{U}}}$ to denote closure under homomorphisms, isomorphisms, subalgebras, products, and ultraproducts, respectively. We will write ${\boldsymbol{\mathit{A}}}\in{{\mathbb{K}}}_{\mathrm{FSI}}$ iff the algebra ${\boldsymbol{\mathit{A}}}\in{{\mathbb{K}}}$ and $\boldsymbol{\mathit{A}}$} is finitely subdirectly irreducible.
@@
  1. 3.
    $\mathbb{K}$}} is the class of all models of some theory, all of whose axioms are equations with the right-hand side equal to 0.
Furthermore, if $\mathbb{K}$}} satisfies any of the equivalent conditions above (i.e., if $\mathbb{K}$}} is a variety), then:
${{\mathbb{K}}}=\mathbf{ISP}({{\mathbb{K}}}_{\mathrm{FSI}}).$
@@
Let ${{\mathbb{K}}}\cup\{{\boldsymbol{\mathit{A}}}\}$ be a class of algebras of the same language $\mathcal{L}$. Let $F$ be a finite subset of $A$ and ${\boldsymbol{\mathit{B}}}\in{{\mathbb{K}}}$. We say that a mapping $f_{F}:F\rightarrow B$ is a partial embedding of the set $F$ from $\boldsymbol{\mathit{A}}$} into $\boldsymbol{\mathit{B}}$} if $f_{F}$ is a one-to-one mapping and for every operation symbol $\lambda\in\mathcal{L}$ of arity $n$, and for all $a_{1},\dots,a_{n}\in F$ such that $\lambda^{{\boldsymbol{\mathit{A}}}}(a_{1},\dots,a_{n})\in F$ we have
@@
We say that $\boldsymbol{\mathit{A}}$} is partially embeddable into the class $\mathbb{K}$}} if for each finite set $F\subseteq A$ there is a ${\boldsymbol{\mathit{B}}}\in{{\mathbb{K}}}$ and there is a partial embedding $f_{F}$ of $F$ into $\boldsymbol{\mathit{B}}$}. If $\boldsymbol{\mathit{A}}$} is partially embeddable into the singleton class $\{{\boldsymbol{\mathit{B}}}\}$, we simply say $\boldsymbol{\mathit{A}}$} is partially embeddable into $\boldsymbol{\mathit{B}}$}.
Theorem 2.3 (Mal’tcev [27]).
@@
    $\mathbb{K}$}} is closed under partial embeddings (i.e. for any algebra $\boldsymbol{\mathit{A}}$} that is partially embeddable into $\mathbb{K}$}} we have ${\boldsymbol{\mathit{A}}}\in{{\mathbb{K}}}$).
@@
An algebra ${\boldsymbol{\mathit{A}}}=\langle{A,+,-,\lor,\land,0}\rangle$ is an Abelian $\ell$-group if $\langle{A,+,-,0}\rangle$ is an Abelian group, $\langle{A,\lor,\land}\rangle$ is a lattice and $\boldsymbol{\mathit{A}}$} satisfies the monotonicity condition: for all $x,y,z\in A$, if $x\leq y$, then $x+z\leq y+z$. A subalgebra of an Abelian $\ell$-group is referred to as an $\ell$-subgroup.
It is well-known (see, e.g. [14, Chapter V]), that the defining conditions of Abelian $\ell$-groups can be expressed by means of equations, so they form a variety which we denote by $\mathbb{AL}$. Also, it is well-known that all Abelian $\ell$-groups are torsion-free.
We will denote by $\boldsymbol{\mathit{Z}}$} and $\boldsymbol{\mathit{R}}$} the $\ell$-groups of integers and reals with the underlying universes $\mathbb{Z}$ and $\mathbb{R}$, taken under their natural order.
For $a\in A$, we define the absolute value $|a|:=a\lor-a$ and we let $n\cdot a:=a+\cdots+a$ denote the $n$-fold sum for every $n\in\mathbb{N}$ (with $0\cdot a:=0$). For $z\in\mathbb{Z}{\setminus}\mathbb{N}$ we set $z\cdot a:=(-z)\cdot(-a)$}.
Definition 2.5.
An $\ell$-subgroup $\boldsymbol{\mathit{B}}$} of an Abelian $\ell$-group $\boldsymbol{\mathit{A}}$} is called convex if it satisfies the following:
$\forall b\in B,\forall a\in A,\quad\big(|a|\leq|b|\implies a\in B\big).$
It is well known that congruences on Abelian $\ell$-groups are in one-to-one correspondence with their convex $\ell$-subgroups (often referred to as $\ell$-ideals), see e.g,̇ [14].
@@
An Abelian $\ell$-group is finitely subdirectly irreducible if and only if it is totally ordered.
@@
All non-trivial totally ordered Abelian $\ell$-groups are universally equivalent. Equivalently, for every non-trivial totally ordered Abelian $\ell$-group $\boldsymbol{\mathit{A}}$} we have $\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{A}}})=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}})$.
From Theorems 2.6 and 2.7 one can easily derive the following:
@@
The following are equivalent for any totally ordered Abelian $\ell$-group $\boldsymbol{\mathit{A}}$}:
@@
Let $\boldsymbol{\mathit{A}}$} be a non-trivial $\ell$-subgroup of $\boldsymbol{\mathit{R}}$}. Then $A$ is a dense subset of $\mathbb{R}$ (with respect to the standard topology on $\mathbb{R}$) or ${\boldsymbol{\mathit{A}}}\cong{\boldsymbol{\mathit{Z}}}$.
@@
Since multiplication by any strictly positive element of $\mathbb{R}$ is an $\ell$-group automorphism on $\boldsymbol{\mathit{R}}$} we can without loss of generality assume $1\in A$. If $A$ also contains some $\xi\in\mathbb{R}\setminus\mathbb{Q}$, we can recursively generate a strictly decreasing sequence of elements of $A$ converging to 0 as follows. We start with $r_{0}=\xi$ and $r_{1}=1$. For $n>0$, we define $k_{n}=\lfloor\frac{r_{n-1}}{r_{n}}\rfloor$ and let $r_{n+1}=r_{n-1}-k_{n}\cdot r_{n}$. Hence, $A$ must be a dense subset of $\mathbb{R}$.
Otherwise, ${\boldsymbol{\mathit{A}}}\in\mathbf{S}({\boldsymbol{\mathit{Q}}})$. By [2, Theorem 2] and [2, Corollary 2] for every non-cyclic ${\boldsymbol{\mathit{A}}}\in\mathbf{S}({\boldsymbol{\mathit{Q}}})$ and for each $0<\epsilon\in\mathbb{R}$ there is an $a\in A$ such that $0<a\leq\epsilon$. Since $\boldsymbol{\mathit{A}}$} is also closed under addition, AA must be a dense subset of $\mathbb{R}$. If $\boldsymbol{\mathit{A}}$} is cyclic, then ${\boldsymbol{\mathit{A}}}\cong{\boldsymbol{\mathit{Z}}}$. ∎
@@
Clearly, the classes $p\mathbb{AL},p\mathbb{AL}^{+},p\mathbb{AL}^{-}$ and $p\mathbb{AL}^{0}$ are varieties. Observe that for an Abelian $\ell$-group $\boldsymbol{\mathit{A}}$} and $a\in A$ we have that $\boldsymbol{\mathit{A}}$} is finitely subdirectly irreducible iff ${\boldsymbol{\mathit{A}}}_{a}$ is finitely subdirectly irreducible. Therefore, using Theorem 2.6 we obtain that a pointed Abelian $\ell$-group is finitely subdirectly irreducible iff it is totally ordered.
As a corollary of Theorems 2.6 and 2.1, together with the fact that every algebra embeds into an ultraproduct of its finitely generated subalgebras (which follows directly from Theorem 2.3), we obtain the following.
Corollary 2.12.
Any subvariety of $p\mathbb{AL}$ is generated by its finitely generated totally ordered members.
Furthermore, the variety $p\mathbb{AL}^{0}$ is clearly term equivalent to $\mathbb{AL}$. We now establish the following result.
@@
    $p\mathbb{AL}_{\mathrm{FSI}}=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$.
@@
    $p\mathbb{AL}_{\mathrm{FSI}}^{+}=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$.
@@
    $p\mathbb{AL}_{\mathrm{FSI}}^{-}=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0})$.
@@
    $p\mathbb{AL}_{\mathrm{FSI}}^{0}=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{0})=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{Z}}}_{0})$.
@@
Let us recall that pointed Abelian $\ell$-groups are finitely subdirectly irreducible iff they are totally ordered.
The inclusions $\supseteq$ are trivial. We prove the remaining inclusions in a convenient order.
@@
    Follows, since the class $p\mathbb{AL}_{\mathrm{FSI}}^{0}$ is term equivalent to $\mathbb{AL}_{\mathrm{FSI}}.$
@@
    Let us fix an arbitrary ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}_{\mathrm{FSI}}$. Using Theorem 2.7 and [9, Lemma 4.3] we obtain ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}(\{{\boldsymbol{\mathit{R}}}_{b}\mid b\in\mathbb{R}\})$. Since for $b\in\mathbb{R}$ we have ${\boldsymbol{\mathit{R}}}_{b}\cong{\boldsymbol{\mathit{R}}}_{-1}$, ${\boldsymbol{\mathit{R}}}_{b}\cong{\boldsymbol{\mathit{R}}}_{0}$, or ${\boldsymbol{\mathit{R}}}_{b}\cong{\boldsymbol{\mathit{R}}}_{1}$ we obtain ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$.
@@
    For ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}_{\mathrm{FSI}}^{0}$ we already know ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$. Therefore, we fix arbitrary ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}_{\mathrm{FSI}}^{+}\setminus p\mathbb{AL}_{\mathrm{FSI}}^{0}$. By the previous point we know that ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1},{\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$. By [3, Theorem 5.6] we have
@@
    Since ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}_{\mathrm{FSI}}^{+}\setminus p\mathbb{AL}_{\mathrm{FSI}}^{0}$, it follows $a>0$. Therefore, ${\boldsymbol{\mathit{A}}}_{a}\notin\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{-1})$ since ${\boldsymbol{\mathit{A}}}_{a}\nvDash\mathtt{f}\leq 0$. Thus ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{R}}}_{0},{\boldsymbol{\mathit{R}}}_{1})$.